Two bodies, A (of mass 1 kg) and B (of mass 3 kg), are dropped from heights of 16m and 25m, respectively. The ratio of the time taken by them to reach the ground is


        (a)    12/5


        (b)    5/12


        (c)    4/5


        (d)    5/4


Grade:10

2 Answers

Jitender Pal
10 years ago

(c)
Let t1 & t2 be time taken to reach the ground then from the formula,
h=   1/2 gt^2,
For first body,   16  =1/2   gt_(1 )^(   2)
For second body,   25 =1/2   gt_2^(   2)
 16/25  =(t_1^(  2))/(t_2^(  2) )   ?  t_1/t_2   =  4/5.

Adarsh
6 years ago
We know v=u+ at ,as it falls downward a=g and u =0V=gt^2g=v/(t^2)For height 16m16/(t1^2)=g -equation 1For height 25m25/(t2^2)=g -equation 2From equation 1 and 2 16/(t1^2)=25/(t2^2)t1/t2=√(16÷25)t1/t2=4/5☺